OpenStudy (smurfy14):

solve: 2^x+2^-x=(65/8)

OpenStudy (anonymous):

Write it as \[2^{-x}(2^{2x}-1)=65/8\]\[2^{-3}(2^{2(3)}+1)=65/8\]\[2^{3}(2^{2(-3)}+1)=65/8\]

OpenStudy (anonymous):

the first should be +1 not -1

OpenStudy (smurfy14):

dont know how you figured out x is 3 but thanks!